# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

### v1

@@ 3.2 Token-Level Evaluation Results
Table 3: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings
Phase Method LR=$3\times 10^{-5}$, $N$=6 LR=$3\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=100
MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
4 A Unified Representational Analysis
@@ 4.1 Representational Analysis Tools
We monitor representational drift with four layer-wise diagnostics: PCA Similarity, PCA Shift, CKA, and the diagonal Fisher Information Matrix (FIM), summarized in Figure 1(b); implementation details (including their precise definitions) are deferred to Appendix A.3.
@@
PCA Similarity & Shift. For each layer $i$, we collect activations $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$. Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,1}^{(*)}$ be he first principal direction and its mean projection for state (){$(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$. The cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$ yields PCA similarity; the signed difference $p_{i,1}^{(*)}-p_{i,1}^{\text{orig}}$ gives PCA shift. Small angles and shifts indicate stable features; otherwise, catastrophic forgetting [36].
@@
Centered Kernel Alignment (CKA). With centered activation matrices $X_{i}^{\text{orig}}$ and $X_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},X_{i}^{(*)})\!\in\![0,1]$; values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
@@
Fisher information. We estimate the diagonal empirical FIM by averaging squared gradients over $\mathcal{X}$. Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning flattens the loss landscape and whether relearning restores parameter importance [14; 11].
@@
We compute all diagnostics not only on the forget set but also on the retain set and on unrelated data, allowing us to distinguish targeted forgetting from broader representational disruption.
@@
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
@@
Figure 2: Layer-wise PCA Similarity for GA on Yi-6B (simple task). (a–c) vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$; (d–f) vary $N\in\{6,50,100\}$ at LR $=3\times 10^{-5}$. Similarity near 1 signals reversible (catastrophic) forgetting; sustained low similarity signals irreversible (catastrophic) forgetting.
@@
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 3: PCA Shift for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Shift magnitude reflects feature displacement: large, unrecovered shifts indicate severe, irreversible (catastrophic) forgetting, while small or fully recovered shifts indicate mild, reversible (catastrophic) forgetting.
4.2 Representational Results
Principal Component Analysis: Similarity and Shift. Figures 2 and 3 show that, in continual unlearning, larger learning rates or more removal requests drive PCA Similarity sharply downward and leave large, unrecovered PCA Shifts—clear evidence of irreversible catastrophic forgetting. Under gentler hyper-parameters, similarity stays high and shifts remain bounded; relearning then realigns both metrics, signalling reversible catastrophic forgetting. Remarkably, the choice of relearning data and analyzed data—forget, retain, or even unrelated—makes little difference: all three restore the feature geometry, implying the knowledge was suppressed rather than erased (Figure 10 and  14). Due to space limitations, complete figures, single-unlearning results, and additional methods appear in Appendix A.5 and Appendix A.6.
Centered Kernel Alignment Analysis. Figure 4 tracks layer-wise CKA. With gentle unlearning, CKA stays near 1 and relearning restores it completely—typical of reversible catastrophic forgetting. Stronger updates or many requests push deep-layer CKA sharply lower, and alignment cannot be fully recovered, signalling irreversible catastrophic forgetting. The choice of relearning or probe data—forget, retain, or unrelated—barely affects this outcome: once relearning begins, latent structure resurfaces regardless of input (Figure 18). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
@@ 4.2 Representational Results
Figure 4: CKA for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. High CKA (near 1) indicates strong subspace alignment and reversible (catastrophic) forgetting, whereas low CKA (near 0) reflects severe representational drift and irreversible (catastrophic) forgetting.
Fisher Information Analysis. Continual unlearning progressively flattens the loss landscape (Figures 5. Higher learning rates or larger request counts push the FIM spectra sharply left, and—at extreme settings—the shift persists after relearning, signalling irreversible catastrophic forgetting. Under gentler hyper-parameters the spectra recentre, indicating reversible catastrophic forgetting. Relearning with forget, retain, or unrelated data realigns the FIM almost equally well, confirming that the lost sensitivity is suppressed rather than erased (Figure 34-38). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) $N=6$, Layer 31
(e) $N=50$, Layer 31
(f) $N=100$, Layer 31
Figure 5: FIM (layer 31) for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Larger leftward shifts indicate a greater flattening of the loss landscape and irreversible (catastrophic) forgetting; spectra that recenter on the original peak denote reversible (catastrophic) forgetting.
4.3 Representational Theoretical Analysis
@@
To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we present a perturbation model linking unlearning updates to structural collapse across layers. Consider an $L$-layer feedforward neural network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$, with activations $\sigma$ and weights ${W_{i}}_{i=1}^{L}$. Unlearning is modeled as layer-wise perturbations $\widetilde{W}_{i}=W_{i}+E_{i}$ with $|E_{i}|=O(\mathrm{LR},N)$, where $\mathrm{LR}$ is the learning rate and $N$ the number of unlearning steps. A Neumann-series expansion yields $\widetilde{f}(x)-f(x)=\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
@@
When small perturbations are confined to a few layers, first-order terms dominate, and the effect is reversible (catastrophic) forgetting. In contrast, comparable perturbations spread across many layers, higher-order terms accumulate, producing irreversible (catastrophic) forgetting.
@@
PCA Similarity. Let $X_{i}$ and $Y_{i}=X_{i}+E^{\prime}_{i}$ be the centered activations at layer $i$ before and after unlearning. By Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$. The layer-averaged PCA similarity is $\bar{S}_{\mathrm{PCA}}\approx 1-O((1/L)\sum_{i}\|E^{\prime}_{i}\|)$.
@@
PCA Shift. Along the first principal component, the activation-centroid shift is $\Delta p_{i}=\mu_{i,1}^{\text{upd}}-\mu_{i,1}^{\text{orig}}=O(\|E_{i}^{\prime}\|)$. Large perturbations $\|E_{i}^{\prime}\|$ spanning many layers cause irreversible representational drift; otherwise, the shifts remain localized and reversible.
CKA. Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ be the perturbed Gram matrix. Then, $\mathrm{CKA}_{i}$ is computed as $1-O(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*})$, which implies that $\bar{C}\approx 1-O((1/L)\sum_{i}\|\Delta K_{i}\|_{*})$.
@@
Token-level metrics, such as accuracy, MIA, and AUC, may report total collapse even when the model’s internal geometry is largely preserved. In reversible forgetting, a few parameter changes (e.g., in output heads or layer norms) can drastically perturb token probabilities while leaving deeper representations intact. For the soft-max output $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2}).$ a small $\delta\theta$ in high-sensitivity regions (near the logits) can dominate the first-order term, producing large drops in accuracy or anomalous AUC scores despite minimal representational drift.
Unlearning on the forget set applies a weight update $E_{i}=LR\times\,\nabla_{W_{i}}L(D_{f}),$ which both removes over-fitting to $\mathcal{D}_{f}$ and accentuates its principal feature subspace. After relearning, parameters may return close to their originals, yet amplified patterns persist, sometimes yielding better performance on an augmented $\mathcal{D}_{f}$ than the baseline. Therefore, unlearning can inadvertently act as a contrastive regularizer, further illustrating the mismatch between surface metrics and feature subspaces.
Summary. Reversible (catastrophic) forgetting occurs when perturbations touch only a few layers; PCA similarity/shift, CKA, and FIM remain near baseline. In contrast, irreversible (catastrophic) forgetting emerges from large, distributed updates that collapse the model’s representational structure.
Unlearning acts as a contrastive perturbation: it removes memorized content yet “reinforces” salient features of the forget set, so subsequent relearning can even outperform the original model on related inputs, revealing its “dual role” in both removal and refinement.
Token-level metrics (e.g., accuracy, MIA) are overly sensitive to small shifts in high-impact parameters and can misclassify the regime. Structural diagnostics, complemented by augmented evaluation, can provide a more reliable assessment of whether forgetting is truly irreversible.
5 Discussion and Takeaways
Beyond theoretical justifications in Section 4, we summarize main empirical and analytical insights.
(1) Single vs. continual unlearning, and the role of GA/RLabel.
Single unlearning rarely produces irreversible collapse: performance is recoverable and representational drift is slight. Continual unlearning, especially with large learning rates, often drives the model into permanent failure: $\sim 100$ sequential requests can push both forget- and retain-set accuracy to near zero. GA and RLabel already over-forget in the single scenario and magnify this damage when applied continually. Adding retain-set terms, as in GA+KL or NPO(+KL) [33; 35; 31], markedly improves stability.
(2) Collapse stems from structural drift, not true erasure.
PCA-Similarity/Shift, CKA, and the FIM consistently expose this breakdown: irreversible collapse coincides with large rotations of principal directions, centroid shifts, and vanishing Fisher mass across many layers. When perturbations remain local (small LR, few unlearning requests), these diagnostics stay near baseline—reversible forgetting. Token-level metrics alone are unreliable: small updates in high-sensitivity parameters (e.g., logits or layer norms) can tank accuracy or inflate MIA AUC while internal geometry is intact.
(3) Unlearning can act as implicit augmentation.
In several continual runs, subsequent relearning on the forget set can often yield higher accuracy than that of the original model. This surprising outcome suggests that unlearning is not merely a memory deletion mechanism but may also serve as a form of implicit contrastive regularization. As detailed in Section 4, unlearning amplifies the feature subspace associated with the forget set, and relearning on augmented inputs can reinforce semantic structure while promoting robustness. This process reorganizes internal representations to better capture generalizable patterns, acting as a form of curriculum learning.
(4) Diagnostics guide irreversible (benign) forgetting.
Tools such as PCA Similarity and Shift, CKA, and Fisher Information reveal not just the presence of structural drift, but also where and how it arises across layers. This enables targeted control: effective unlearning can be guided toward perturbing only those parameters responsible for the forget set, while preserving the representation structure on the retain set and unrelated data. This opens a path to targeted, irreversible forgetting, a permanent and isolated removal of information without collateral collapse, offering actionable insights for building safer unlearning algorithms.

### v2

@@ 3.2 Evaluation Results
Continual Unlearning. Post-relearning analysis (Tables 3, 6, and 7) reveals two forms of reversible forgetting. In reversible, catastrophic forgetting, both utility (e.g., F.Acc, R.Acc) and privacy metrics drop sharply during unlearning but are fully restored after relearning. This is observed in GA and RLabel with moderate hyperparameters. Besides, reversible, non-catastrophic forgetting entails only a mild, easily recoverable performance degradation, as seen with NPO at $\mathrm{LR}=3\times 10^{-5},\,N=6$.
Conversely, irreversible, catastrophic forgetting occurs when relearning fails to restore utility, leaving F.Acc and R.Acc low despite partial MIA recovery. This pattern is common for GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5},N=100$), where cumulative updates lead to irreversible representational collapse. The MIA AUC metric behaves erratically in this regime: it may fall below 50% during unlearning but misleadingly rebounds to high values after relearning, even after the model’s capabilities have been permanently lost. These empirical results on single and continual unlearning are consistent with the theoretical framework in Section 5, which shows that small perturbations to the model weights can trigger disproportionately large drops in accuracy.
4 Representation-level Evaluation
@@ 4.1 Representational Analysis Tools
To analyze representational drift, we employ four hidden state diagnostics, as summarized in Figure 1(b). Their precise definitions and implementation details are deferred to Appendix A.3.
@@
PCA Similarity, Shift, and Mean Distance. For each layer $i$, we collect activation matrices $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$ for the original, unlearned, and relearned models, respectively. Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,12}^{(*)}$ be the first principal direction and its mean projection for state $(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$. PCA Similarity is the cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$, while PCA Shift is the signed difference $p_{i,12}^{(*)}$. Small values for these metrics indicate stable features, whereas large, unrecovered shifts signify irreversible changes [45]. We also introduce the mean PCA distance, the average Euclidean distance on $p_{i,12}^{(*)}$ across layers, to provide a single scalar measure of representation drift.
@@
Centered Kernel Alignment (CKA). Given centered activation matrices $X_{i}^{\text{orig}}$ and $Y_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},Y_{i}^{(*)})\!\in\![0,1]$. Values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
@@
Fisher information (FIM). We estimate the diagonal of the empirical FIM by averaging squared gradients over the probe set $\mathcal{X}$. Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning alters the loss landscape and whether relearning restores parameter importance [15; 10].
@@
All diagnostics are computed not only on the forget set but also on the retain set and unrelated data to distinguish targeted unlearning from general representational degradation.
4.2 Representational Results
(a) Simple (Single Unlearning)
(b) Simple (Single Unlearning)
(c) Simple (Single Unlearning)
(d) Simple (Single Unlearning)
Figure 2: Single unlearning analysis on Yi-6B with GA under a simple task. In reversible non-catastrophic forgetting, PCA Similarity, PCA Shift, CKA, and FIM across layers show only minor changes with slight accuracy drops. Input queries are drawn from the forget set.
4.2.1 Single Unlearning
Figure 2 demonstrates feature-level changes under single unlearning. (a) PCA Similarity remains near 1.0 across all layers, with minor, reversible dips, indicating that dominant activation directions are preserved. Slight dips in shallow and final layers are rapidly restored after relearning, suggesting minimal and reversible drift. (b) PCA shifts are minimal, and relearned representations closely realign with the original. (c) CKA values are nearly 1.0 for all model states, confirming that subspace structures remain intact. (d) FIM spectra show only mild, temporary shifts that are fully restored after relearning. These results, combined with the task-level evaluation in Section 3.2, demonstrate that single unlearning induces reversible, non-catastrophic forgetting. This highlights the limitation of classic (task-level) metrics, which fail to capture the superficial nature of the forgetting.
4.2.2 Continual Unlearning
As shown in Figures 3 and 5, PCA Similarity and Shift offer complementary views of representational change: similarity reflects global alignment, and shift is more sensitive to local variations. Relying on PCA similarity alone can obscure subtle effects; employing both avoids overlooking fine-grained distinctions, enabling a more comprehensive assessment. Higher learning rates or more requests cause sharp drops in similarity and large, unrecovered shifts, which is characteristic of irreversible catastrophic forgetting. In contrast, milder hyperparameters lead to high similarity and bounded shifts that are restored after relearning, consistent with reversible, catastrophic forgetting. This pattern is consistent across probe sets from forget set, retain set, and unrelated data (Figures 10 and 14).
Figure 4 integrates CKA (top) and FIM (bottom) analyses. CKA reveals that mild unlearning maintains stable alignment that recovers post-relearning, while aggressive unlearning causes irreversible degradation. The FIM spectra complement this by showing that continual unlearning flattens the loss landscape. Extreme hyperparameters induce a permanent leftward shift in sensitivity distributions, whereas moderate settings permit recovery. Together, these diagnostics suggest that observed performance loss is often due to temporary suppression rather than permanent erasure of knowledge. For conciseness, we present results on forget-set queries in the main text; retain-set queries, which yield similar conclusions under catastrophic forgetting, are in Appendix A.5 (e.g., Figures 12 and 10).
Mean PCA Distance Analysis. To quantify representation-level drift with a single metric, we use the mean PCA distance. To assess the metric’s sensitivity, we compute its mean and standard deviation across four random seeds, as well as shuffling the order of unlearning requests. As shown in Table 4, higher learning rates consistently increase the mean PCA distance for both unlearned and relearned models At a low learning rate (e.g., $3\times 10^{-6}$), the mean and variance of the distance remain low after relearning, indicating stable and reproducible recovery. In contrast, a high learning rate (e.g., $3\times 10^{-5}$) sharply increases both values, reflecting greater variability and incomplete recovery.
Combining these empirical findings with the task-level results in Section 3, we further connect to the theoretical analysis in Section 5, whose conclusions show that small weight perturbations induce only limited changes in the feature space, thereby enabling recovery, whereas larger perturbations accumulate into substantial representational drift that underlies irreversibility.
@@ 4.2.2 Continual Unlearning
(d) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
@@
Figure 3: Layer-wise PCA Similarity and Shift for GA on Yi-6B (simple task). Vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$. Sustained low similarity or large shifts signal severe, irreversible catastrophic forgetting, whereas partial similarity or small shifts indicate mild, reversible catastrophic forgetting. Input queries are drawn from the forget set.
@@
[…]
@@
Figure 4: CKA for GA on Yi-6B, simple task. Vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$. High CKA ( 1) and concentrated FIM spectra indicates reversible catastrophic forgetting, while persistently low CKA and large-shifted, flattened spectra denote severe representational drift and irreversible catastrophic forgetting. Input queries are drawn from the forget set.
Table 4: Yi-6B (GA): Mean $\pm$ standard deviation of PCA distance on forget and retain sets across varying learning rates and random seeds; the number of unlearning requests is fixed.
Model Learning Rate Phase Seed Mean PCA distance (forget set) Mean PCA distance (retain set)
[…]
$\mathbf{3\times 10^{-6}}$ Relearn 12, 22, 32, 42 $2.11\pm 1.42$ $1.64\pm 1.12$
$\mathbf{5\times 10^{-6}}$ Unlearn 12, 22, 32, 42 $11.52\pm 6.19$ $8.79\pm 5.20$
$\mathbf{5\times 10^{-6}}$ Relearn 12, 22, 32, 42 $1.37\pm 0.74$ $1.05\pm 0.58$
$\mathbf{3\times 10^{-5}}$ Unlearn 12, 22, 32, 42 $133.20\pm 45.81$ $121.45\pm 38.60$
$\mathbf{3\times 10^{-5}}$ Relearn 12, 22, 32, 42 $104.58\pm 39.70$ $95.34\pm 32.40$
5 Theoretical Analysis
5.1 A Perturbation Model of Unlearning
@@
To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we introduce a perturbation model that links unlearning updates to representational changes across layers. Consider an $L$-layer feedforward network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$ with activation $\sigma$ and weights ${W_{i}}_{i=1}^{L}$. We model unlearning as a layer-wise perturbation $\widetilde{W}_{i}=W_{i}+E_{i}$, where the magnitude of the error term, $|E_{i}|=O(\mathrm{LR},N)$, scales with the learning rate ($\mathrm{LR}$) and the number of unlearning requests $N$. A Neumann-series expansion of the network’s output shows that the total change $\widetilde{f}(x)-f(x)$ is defined as $\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
@@
When perturbations are small and localized to a few layers, first-order terms dominate, leading to reversible (catastrophic) forgetting. In contrast, when comparable perturbations are distributed across many layers, higher-order interaction terms accumulate, causing structural degradation that results in irreversible (catastrophic) forgetting. We can formalize the impact on our diagnostic tools:
@@
PCA Similarity. Let $X_{i}$ and $Y_{i}=X_{i}+E^{\prime}_{i}$ be the centered activations at layer $i$ before and after unlearning. By the Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$. The layer-averaged PCA similarity is $\bar{S}_{\mathrm{PCA}}\approx 1-O((1/L)\sum_{i}\|E^{\prime}_{i}\|)$.
@@
PCA Shift. Along the first principal component, the activation-centroid shift is expressed as $p_{i,12}=O(\|E_{i}^{\prime}\|)$. Large perturbations $\|E_{i}^{\prime}\|$ propagating across multiple layers lead to irreversible representational drift, whereas smaller perturbations remain localized and thus reversible.
CKA. Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ denote the perturbed Gram matrix at layer $i$. The corresponding CKA score is computed as $\mathrm{CKA}_{i}=1-O\!\left(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*}\right)$. Averaging across layers yields $\bar{C}\approx 1-O\!\left(\tfrac{1}{L}\sum_{i}\|\Delta K_{i}\|_{*}\right)$, where $\bar{C}$ denotes the layer-averaged CKA.
@@
Table 5: Yi-6B (GA+GD+WAGLE) performance under different relearning settings. F.Acc/R.Acc are forget/retain accuracy, with mean PCA distances on the forget and retain sets.
Phase F.Acc R.Acc Mean PCA distance (forget set) Mean PCA distance (retain set)
Original model 78.9 65.5 0 0
[…]
Unlearn 37.8 55.9 11.84 6.28
Relearn 46.9 58.3 9.00 5.91
[…]
Unlearn 27.8 51.4 26.02 8.37
Relearn 31.5 53.5 24.56 8.12
5.2 Bridging Representational Drift and Task-Level Metrics
Classic (task-level) metrics can be misleading. They are highly sensitive to small weights’ changes, particularly in the final layers, which can cause large shifts in output probabilities without altering the model’s deeper representations. For a softmax output, a small perturbation $\delta\theta$ to the model parameters yields a large change in log-probability: $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2})$. A minor update to the logits can dominate this first-order term, causing a sharp drop in accuracy that suggests catastrophic forgetting, even if the underlying geometry is preserved.
This aligns our theoretical model with the empirical findings in Sections 3 and 4.2. When $\mathrm{LR}$ or $N$ is small, changes are confined to first-order effects, feature spaces remain intact, and forgetting is reversible. When $\mathrm{LR}$ or $N$ is large, higher-order perturbations accumulate across layers, making recovery impossible and leading to irreversible forgetting. Figure 3 illustrates such a transition.
Interestingly, relearning can sometimes lead to performance that exceeds the original model’s accuracy on the forget set (Table 3). This suggests that unlearning can act as a form of contrastive regularization, reinforcing salient features related to the forgotten data, which a brief relearning can then exploit.
5.3 Probing the Limits of Irreversibility
In our primary experiments, we did not observe irreversible non-catastrophic forgetting; even a small fraction (e.g., $10\%$) of the forget set was sufficient to restore performance. To explore this regime, we conducted extra experiments with more constrained relearning conditions. We used the GA+GD+WAGLE method [13], which selectively updates influential parameters, and limited the relearning data to either (i) $50\%$ of the retain set or (ii) an equal-sized, unrelated dataset (Table 5).
Under these conditions, the method exhibited seemingly irreversible, non-catastrophic forgetting. The forget set showed large, unrecoverable PCA distances, while the retain set experienced only modest, partially recoverable degradation. This demonstrates that achieving the ideal of targeted, permanent unlearning without collateral damage remains an open challenge. Defining precise thresholds to distinguish the forgetting regimes (reversible vs. irreversible and catastrophic vs. non-catastrophic) is non-trivial, as they depend on the unlearning method, task complexity, and other factors.
Our representational diagnostics reveal structural drift on each layer, pointing toward avenues for more robust and effective unlearning. A promising future direction is the design of new loss functions, such as a forget component penalization term to dismantle target representations and a retain component regularization term to preserve the geometry of desired representations.